A gas occupies a volume of 64.2L at a given temperature and pressure. What will be the volume of this gas if we reduce the absolute temperature by half that quadruple the pressure and a third of the particles escape from the container
Answer #1

Tf = Ti/2

Pf = 4*Pi

nf = 2/3 ni

Use:

Pi*Vi/ni*Ti = Pf*Vf/nf*Tf

Pi*64.2/ni*Ti = 4Pi*Vf/ (2/3ni*Ti/2)

64.2 = 4Vf/ (2/6)

64.2 = 4Vf*6/2

64.2 = 12Vf

Vf = 5.35 L

Answer: 5.35 L